At noon a train leaves Washington, D.C., headed for Charleston, South Carolina, a distance of 500 miles. the train travels at a speed of 60 mph. At 1 p.m. a second train leaves charleston headed for washington d.c., traveling at 50 mph. how long after the train leaves charleston will the twon trains pass each other?
Answer by mananth(16949)

At noon a train leaves Washington, D.C., headed for Charleston, South Carolina, a distance of 500 miles. the train travels at a speed of 60 mph. At 1 p.m. a second train leaves charleston headed for washington d.c., traveling at 50 mph. how long after the train leaves charleston will the twon trains pass each other?
...
12 noon train A - speed = 60 mph
1 pm train B speed 50 mph
...
when train B starts train A has traveled 60 miles.
Balance distance between two trains = 440 miles.
...
they are travelling towards each other.
so speed = 60+50=110 mph.
distance 440 miles
time = 440/110
time = 4 hours
they will meet 4 hours after train B starts
...
m.ananth@hotmail.ca